# ImageSet2Text: Describing Sets of Images through Text

### v1

@@ Results.
4.3 Readability and Overall Quality
The previous two experiments provided standardized metrics of the accuracy and completeness of the descriptions in two downstream tasks. However, they did not directly assess the quality of the descriptions themselves. To address this gap, we conducted a user study.
Methodology.
We randomly selected $60$ image sets from PIS, $20$ from each difficulty category, and generated descriptions using ImageSet2Text. From each image set, we randomly selected $16$ images, and displayed them in a $4\times 4$ grid alongside their corresponding descriptions, as depicted in App. D. Participants were asked to answer $5$ five-point Likert-scale questions (where $1$ corresponds to the lowest score) to evaluate, for each description: (1) its clarity; (2) its accuracy; (3) its level of detail; (4) how natural its flow is; and (5) the overall satisfaction of the participant with the description.
We compared the ImageSet2Text descriptions with control descriptions. Since no existing baseline generates descriptions of large image sets, we created $10$ control descriptions of three different types: Control Accuracy ($3$ descriptions), which consist of well-written but highly inaccurate descriptions. For example, if the images depict cars, the control description could describe dogs; Control Detail ($3$ descriptions), which consist of descriptions that refer to the right visual content in the images, but lack detail and are overly vague; and Control Clarity/Flow ($4$ descriptions), which accurately describe the images with sufficient amount of detail but lack a natural flow. Control descriptions were generated with ChatGPT and some examples are provided in App. D.
Participants.
A total of $233$ people were reached through the study, each of whom was asked to evaluate $7$ descriptions, of which $6$ were descriptions by ImageSet2Text and $1$ was a control description. The study included $2$ attention checks to filter out participants who did not pay enough attention to the task. In total, 198 participants successfully completed the study. Each set-description pair was evaluated by a minimum of 16 and a maximum of 22 participants (avg. 19.8). The participants were recruited through Prolific, [^3] Prolific, https://www.prolific.com/, Last Access: 7th of March 2025. after qualifying for the study. [^4] To qualify for the study, the participants had to be adults (at least 18 years old), native English speakers, and could not have visual impairments or reading comprehension difficulties. The overall task took around $8$ minutes to complete and the successful participants were compensated with an hourly rate of \$$\$12$. The data collection was fully anonymized, and no personal information was collected.
Results.
The results are depicted in Fig. 3. The descriptions generated by ImageSet2Text were rated positively, both generally and relative to the reference levels of the control descriptions: clarity ($\mu=4.29$ vs $\mu=2.80$ for the control), accuracy ($\mu=3.76$ vs $\mu=1.43$ for the control), level of detail ($\mu=4.06$ vs $\mu=2.96$ for the control) and flow ($\mu=3.96$ vs $\mu=2.07$ for the control). Based on these human evaluations, we conclude that the descriptions created by ImageSet2Text are clear, readable and with an overall good quality.
Figure 3: Boxplots of the human evaluation of the descriptions generated by ImageSet2Text and the control descriptions. The values of the control descriptions correspond to those designed to assess clarity, accuracy, detail and flow, whereas the ratings provided to all the control descriptions together are used to assess overall an satisfaction. The magenta bar corresponds to the median while the magenta dot represents the mean, numerically reported in the figure.
5 Ablation Study
The integration of structured data representations, such as graphs or dependency parsing, into data-driven AI systems is a key area of ongoing research, particularly in the context of the evolving debate between symbolic vs data-driven AI [42, 21]. We conducted an ablation study to examine the role of structured representations in the predominantly data-driven pipeline of ImageSet2Text.
The ablation study considers four versions of ImageSet2Text, each progressively introducing structured information into different components:
@@ 5 Ablation Study
    relies only on LLMs and CVLs, without using any graph representation at any stage of the pipeline. At each iteration, the hypothesis $h$, the set of more general hypotheses $\mathcal{H}$, and the supporting $\mathcal{H}_{i}^{+}$ and contradicting $\mathcal{H}_{i}^{-}$ alternatives are generated by prompting the LLM. Rather than storing information in an intermediate graph representation $\mathcal{G}_{\tau}$, the extracted insights from each round are directly used to iteratively refine a textual description $d_{\tau}$;
@@
    introduces the knowledge graph to generate the set $\mathcal{H}$ of more general hypotheses, and the supporting $\mathcal{H}_{i}^{+}$ and contradicting $\mathcal{H}_{i}^{-}$ alternatives. However, no graph representation is kept in memory and the textual description is updated at every iteration;
@@
    summarizes the dataset $\mathcal{D}$ not anymore through an iterative textual description but through the graph representation $g_{\tau}$, kept in memory and used to generate a more concise description at the end of the iterative process;
@@
    utilizes the LLM to summarize the answers from the VQA into a sentence, which is then processed through dependency parsing to identify entities and relations and generate the hypothesis $h$.
Table 4: ImageSet2Text ablation study with incremental structured knowledge representation on a subset of PIS dataset [18].
graph to create graph stored dependency
LLM hypotheses in memory parsing Acc@1 Acc@5
@@
We compare the completeness performance as in Sec. 4.2 on a random subset of $15$ ($5$ easy, $5$ medium, and $5$ hard) image set pairs, reporting both accuracies at $1$ and at $5$. In addition, a manual assessment of the quality of the descriptions is conducted by two of the authors. As reported in Tab. 4, the progressive integration of structured information improves the performance up to v3, while there is a drop in performance in v4. This result is confirmed in the manual assessment of the descriptions, where the quality of the descriptions generated by v3 was clearly superior to that of the other versions. Being the best-performing version, v3 is the version presented and evaluated in the paper.
Next, we provide a case-by-case comparison. The generation of the sets $\mathcal{H}$, $\mathcal{H}_{i}^{+}$, and $\mathcal{H}_{i}^{-}$ in v2 is consistently more streamlined compared to v1 that only relies on the LLM which is subject to hallucinations for this task. However, the descriptions generated by v2 and v1 have sometimes a broken flow and are hard to follow. This problem is solved in v3 by introducing the final description generated directly from the graph. Finally, while v4 includes dependency parsing to generate the candidate predicates of the nodes, such a feature does not seem necessary as this specific task already benefits from the rich contextual embedding space of the LLM.
From this ablation study, we conclude that the best-performing version, v3, is the version that best leverages the advantages of both structured and data-centric AI.
6 Limitations and Future Work
While the experiments and ablation study confirm the strengths of ImageSet2Text, they also reveal areas for improvement especially in the hypothesis verification phase. These issues arise from limitations in the CVL embeddings, which enable scalability to large image collections but at times lead to false positives/negatives, and in the use of WordNet to create contradicting hypotheses because sibling nodes may not be mutually exclusive. Additionally, semantically different concepts in WordnNet which are visually indistinguishable in the CLIP embedding space, such as “tea” vs. “tea-like”, remain a challenge. We leave to future work addressing these limitations. Furthermore, ImageSet2Text uses GPT-4o-mini as the LLM, which is a proprietary model. In future work, we plan to explore open-source alternatives for accessibility and ethical considerations.
@@ 6 Limitations and Future Work
The high readability of the descriptions generated by ImageSet2Text suggests broader applications, including dataset exploration under user-guidance, explainable AI and the identification of potential biases in image sets. In addition, the experiment on GroupWikiArt highlights promising possibilities in the field of cultural analytics [40]. Exploring the applicability of ImageSet2Text in real-world scenarios, we have established a collaboration with Fundación ONCE [^5] Fundación ONCE, https://www.fundaciononce.es/en, Last Access: 19th of March 2025., a Spanish NGO devoted to improving universal accessibility, to gather their insights on the potentialities of developing use cases of ImageSet2Text for improving the quality of life of visually impaired individuals. As authors of this paper, we believe that community insights should be taken into account early in the design process [13]. Details of these discussions are provided in App. E.
@@ 7 Conclusion
We have presented ImageSet2Text, a system to automatically generate natural language descriptions of image sets, a novel task in the literature. To assess the accuracy of these descriptions, we have conducted a large-scale group image captioning experiment and released two benchmark datasets: GroupConceptualCaptions and GroupWikiArt. We have further demonstrated their completeness through strong performance in the Set Difference Captioning task. Additionally, a human evaluation by means of a user study confirmed the readability and overall quality of the generated descriptions. Since ImageSet2Text leverages both structured and data-centric approaches, we have performed an ablation study that offers insights into the value of integrating these two paradigms.
@@ Acknowledgements
We are grateful to Fundación ONCE for the fruitful collaboration and their willingness to collaborate on our research.
PR and NO are supported by a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Convenio Singular signed with Generalitat Valenciana, Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación). PR is also supported by a grant by Fundación Banc Sabadell.
A part of this work was performed while PR was an academic guest at the University of Osaka, in the D3 center. NG is partly supported by JSPS KAKENHI No. JP22K12091.
KS is part of the ELLIS Unit Linz, the LIT AI Lab and the Institute for Machine Learning at Johannes Kepler University Linz, which are supported by the Federal State Upper Austria. This research was funded in part by the Austrian Science Fund (FWF) [10.55776/COE12]. We thank the projects INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
We thank Erik Derner, Korbinian Pöppel, Fabian Paischer, and Aditya Gulati for helpful discussions on different aspects of this work.
Supplementary Material
A ImageSet2Text: Implementation Details
@@
To effectively prompt GPT-4o-mini, we designed system prompts that outline the model’s role and guide its reasoning process. Additionally, we utilized structured outputs to ensure consistency in responses. For example, we reference the “next question” prompt used in ImageSet2Text, provided in Fig. 4, to explain the rationale behind our prompt engineering. This is a key component in our design, as it provides a structured method for generating follow-up questions to refine image set descriptions.
The system prompt defines GPT-4o-mini’s role as an expert assisting in the enrichment of textual descriptions for a large image set. Specifically, the model is given:
@@
KEY_POINT: The main element within the description that needs additional details.
ATTRIBUTE: The specific property of the KEY_POINT to investigate further.
@@
This structured approach overall improves the quality of GPT-4o-mini. The other prompts used for the LLM in ImageSet2Text follow the same criteria.
NEXT_QUESTION = """You are an expert in {expertise} assisting a client in enriching the textual description of a large IMAGE_SET in their possession.
Using JSON, you will be provided with a question expressed as a connection within a knowledge graph built for IMAGE_SET (QUESTION_BRANCH), a KEY_POINT that lacks detailed explanation in DESCRIPTION_CURRENT, an ATTRIBUTE which is the specific property of the KEY_POINT to investigate further, and a list of textual questions which have already been asked about that KEY_POINT (LOG).
An example of a client query might be the following:
{jsonScheme_input}
Based on this input, translate the QUESTION_BRANCH into two textual concise questions (QUESTION_EXPERT and QUESTION_VQA) to investigate the ATTRIBUTE further.
Provide the answer in JSON format. For example, the output from an expert in couples would be:
{jsonScheme_output}
In the output, it is expected that:
- The field QUESTION_EXPERT is a string containing an expert-level query that specifically addresses the selected ATTRIBUTE, intended to enrich the understanding of the IMAGE_SET.
- The field QUESTION_VQA is a string containing the translation of QUESTION_EXPERT into an image-focused question that meets the following criteria:
1. QUESTION_VQA must be direct, simple, clear, and must refer to visible aspects within the image.
2. QUESTION_VQA should be structured to ask about details that can be observed in one randomly selected image from the IMAGE_SET, referring to "the image" directly, like "What is in the image?".
3. Avoid questions that require abstract reasoning, cultural knowledge, or expertise beyond what can be visually interpreted.
4. Avoid yes/no questions and focus instead on generating descriptive responses.
5. Ensure that QUESTION_VQA aligns with the capabilities of a VQA system, such as object recognition, spatial relationships, counting, and scene understanding.
6. QUESTION_VQA must be designed to elicit new insights without overlapping with previous responses.
@@
def next_question(expertise):
expertise = expertise_to_string(expertise)
jsonScheme_input = """{
␣␣␣␣␣␣␣␣"QUESTION_BRANCH":"image.wedding.couple.body language?",
␣␣␣␣␣␣␣␣"KEY_POINT":"couple",
␣␣␣␣␣␣␣␣"ATTRIBUTE":"body language: communication via the movements or attitudes of the body",
␣␣␣␣␣␣␣␣"log":[
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣"ATTRIBUTE":"gender composition: the properties that distinguish organisms on the basis of their reproductive roles",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣"QUESTION_EXPERT":"What is the gender composition of the couple portrayed in the image?",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣"QUESTION_VQA":"How many men and women are visible in the image?"
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣"ATTRIBUTE":"attire: clothing of a distinctive style or for a particular occasion",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣"QUESTION_EXPERT":"What is the attire of the couple portrayed in the image?",
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣"QUESTION_VQA":"What type of clothing are the individuals in the image wearing?"
jsonScheme_output = """{
␣␣␣␣␣␣␣␣"QUESTION_EXPERT":"How does the couple’sbodylanguageappearintheweddingimage?",
␣␣␣␣␣␣␣␣"QUESTION_VQA":"Whatarethebodypositionsormovementsofthespousesintheimage?"
␣␣␣␣returnNEXT_QUESTION.format(
␣␣␣␣␣␣␣␣expertise=expertise,
␣␣␣␣␣␣␣␣jsonScheme_input=jsonScheme_input,
␣␣␣␣␣␣␣␣jsonScheme_output=jsonScheme_output
@@
B Accuracy Evaluation
In this section, we provide additional information regarding the accuracy evaluation. Specifically, we provide details about the dataset generation process, how to create a caption with ImageSet2Text, details on how the baselines have been set up and the full results of the experiments.
B.1 Dataset Composition and Creation
Given the absence of public benchmarks for evaluating ImageSet2Text in the group image captioning task, we constructed datasets. Specifically, the GroupConceptualCaptions dataset based on the ConceptualCaptions dataset [55] and the GroupWikiArt dataset based on the WikiArt dataset [57]. Group sizes for these datasets are depicted in Fig. 5. The minimal size is 50, and many groups are smaller than 1000 samples, yet there are also sets in the range of 3000-4000 images.

### v2

@@ Results
4.3 User Satisfaction
While the previous experiments measure accuracy and completeness, they do not assess the overall quality of the descriptions, which we evaluate in a user study conducted on a random subset of the PIS dataset with $233$ users.
Methodology
We sample 60 PIS image sets (20 per difficulty level) and, for each, we show participants 16 images in a $4\times 4$ grid alongside a description (see App. D). Participants are asked to rate the clarity, accuracy, detail, flow, and overall satisfaction of the descriptions on a five $5$-point Likert-scale. To ease the interpretation of the user feedback and given that there is a lack of alternative methods to create descriptions of large image sets, we generate $10$ control descriptions using ChatGPT (see App. D for examples), divided into three categories:
  • Control accuracy (3 descriptions): well-written but intentionally inaccurate descriptions.
  • Control detail (3 descriptions): descriptions that reference the correct visual content but lack details.
  • Control clarity and flow (4 descriptions): factually correct, detailed but with low coherence descriptions.
Note that comparing against the captions described in Section 4.1 would be inappropriate due to length and detail mismatches that could bias results (grice1975logic; kahneman2011thinking). Also note that the control descriptions are not performance baselines: they isolate specific qualities, allowing for controlled comparisons in alignment with human-centered evaluation best practices (rogers2023interaction).
A total of $233$ qualifying participants55518+ native English-speakers w/o visual/reading impairments. were recruited via Prolific, [^6] https://www.prolific.com/ (accessed 2025/03/07). each evaluating $7$ descriptions ($6$ ImageSet2Text, $1$ control) and performing $2$ attention tests. $198$ participants successfully completed the study, yielding 162216-22 evaluations per description. The average task completion time was $8$ minutes, compensated at \$12/hour. The process was fully anonymized, and no personal information was collected.
Results
Results are shown in Fig. 5. ImageSet2Text descriptions are consistently rated favorably, not only in absolute terms but also when compared to the reference levels established by the control descriptions: clarity ($\mu=4.29$ vs control $\mu=2.80$), accuracy ($\mu=3.76$ vs control $\mu=1.43$), detail ($\mu=4.06$ vs control $\mu=2.96$) and flow ($\mu=3.96$ vs control $\mu=2.07$). These differences are statistically significant (t-test, all $p\ll 10^{-5}$).
Figure 5: User study results. For control, values are of those designed to assess clarity, accuracy, detail, and flow, respectively, whereas overall and satisfaction are averaged across all control descriptions. Purple bars indicate medians, dots show means with the actual values reported in the figure.
5 Method’s behavior
We report analyses on ImageSet2Text’s behavior, considering an ablation study, scalability estimate and failure cases analysis (Fig. 2) and discussing potential applications.
Ablation study
Inspired by ongoing research that combines symbolic and data-driven AI (marcus2018; guo2024), we conducted an ablation study with four versions of ImageSet2Text, each progressively integrating more structured information:
@@ 5 Method’s behavior
    relies only on LLMs to generate hypothesis $h_{0}$, the set of general hypotheses $\mathcal{H}$, and the supporting $\mathcal{H}_{i}^{+}$ and contradicting $\mathcal{H}_{i}^{-}$ alternatives. The extracted insights are directly used to refine the description iteratively;
@@
    introduces the external lexical graph $\mathcal{G}_{l}$ to generate the sets $\mathcal{H}$, $\mathcal{H}_{i}^{+}$ and $\mathcal{H}_{i}^{-}$. No concept graph is kept in memory and the textual description is updated at every iteration;
@@
    introduces the iterative concept graph $\mathcal{G}_{c}^{\tau}$, with the final description being generated from $\mathcal{G}_{c}^{T}$. This is the version of ImageSet2Text introduced in this paper;
@@
    decomposes hypothesis formulation into two steps: (1) the LLM to summarize the VQA answers $\mathcal{A}$ into a sentence, and (2) POS tagging and dependency parsing to extract the object $o_{0}$ and its candidate predicates.
Table 2: Ablation study with incremental structured knowledge representation on a subset of the PIS dataset (* indicates the version corresponding to ImageSet2Text).
$\mathcal{G}_{l}$ $\mathcal{G}_{c}^{\tau}$  dependency parsing Acc@1 Acc@5
@@
We report completeness (as in Section 4.2) on a random subset of $15$ image set pairs ($5$ easy, $5$ medium, and $5$ hard). Table 2 shows that the progressive integration of structured information improves the performance up to v3, followed by a decline in v4. This result is confirmed by manual assessment, where v3 produced the highest quality descriptions. The generation of $\mathcal{H}$, $\mathcal{H}_{i}^{+}$, and $\mathcal{H}_{i}^{-}$ in v2 is more effective compared to v1, which is subject to hallucinations due to relying only on the LLM. However, both v2 and v1 occasionally produce descriptions with a broken flow. This limitation is addressed in v3 by directly generating the final description from the concept graph. POS tagging and dependency parsing in v4 are error-prone and hard to adapt, making them less reliable than the LLM alone for hypothesis formulation. In conclusion, the best-performing version, v3, is the version that best leverages the advantages of both symbolic and data-centric AI.
Scalability
ImageSet2Text scales efficiently to large image sets of size $N$. Hypotheses are generated from a small random sample $M\ll N$, keeping computational cost independent of $N$ (with $M$ depending on set heterogeneity rather than size). For verification, $1280$-dimensional embeddings (i.e., $d=1280$) for all $N$ images are precomputed once on the CVL model at $\approx$ $12$ images/s on an NVIDIA RTX 3090, requiring $T_{embed}(N)=N/12$ seconds. [^7] https://gist.github.com/TACIXAT/ecd4f636bf6af28cb69d641e29d7b362 (accessed 2025/07/28) The kNN step, repeated over $T$ iterations, compares $N$ images with $S+C$ example points, taking $T_{iter}(N,|\mathcal{H}_{i}^{+}|,|\mathcal{H}_{i}^{-}|)=N\cdot(|\mathcal{H}_{i}^{+}|+|\mathcal{H}_{i}^{-}|)\cdot d\cdot 2/\mathrm{FLOPs}$ seconds, where $\mathrm{FLOPs}$ denotes the GPU’s floating-point throughput. The factor of 2 accounts for one multiplication and one addition per dimension when computing cosine distance on L2-normalized embeddings. For example, using FP32 precision on an NVIDIA RTX 3090 (35.58 TFLOPs), with $1$ million images and $|\mathcal{H}_{i}^{+}|+|\mathcal{H}_{i}^{-}|=1000$, each kNN iteration takes <0.1<0.1 s, which is ImageSet2Text’s only $N$-dependent step. Across our experiments, $T$ ranges from $10$ to $30$, independently of $N$. In addition, avg graphs #\#nodes (depth) are: 10.84 (3.28) in GroupConceptualCaptions, 9.83 (3.13) in GroupWikiArt; in PIS: 10.84 (3.31) easy, 10.48 (3.48) medium, 10.37 (3.57) hard, indicating fewer nodes but similar/higher depth due to higher visual diversity.
Figure 6: Illustrative examples of failure cases: (1) sibling nodes that are not mutually exclusive (pachyderm vs. proboscidean), causing incorrect formulation of $\mathcal{H}_{i}^{-}$; and (2) sibling nodes that cannot be visually distinguished (peach vs. nectarine), causing kNN misclassifications.
Failure cases
We identify two main failure types (see Fig. 6) due to the integration of WordNet and CLIP: (1) sibling nodes that are not mutually exclusive cause $\mathcal{H}_{i}^{-}$ elements to be supportive (in 1.49% of cases); (2) sibling nodes that are hard to distinguish visually cause kNN misclassifications (in around 2% of cases). While these issues affect the verification process, they are sufficiently rare to have minimal overall impact. Details on how we estimate their incidence are provided in App. E, along with other limitations inherited from the modules constituting ImageSet2Text.
Potential Applications
@@
The high readability of the descriptions generated by ImageSet2Text suggests broader applications beyond the already demonstrated possibilities of the SDC task (visdiff) (Section 4.2), such as dataset exploration under user-guidance, explainable AI, cultural analytics, and the identification of potential biases in image sets. Additionally, we initiated a collaboration with Fundación ONCE [^8] https://www.fundaciononce.es/en (accessed 2025/11/11) to collect feedback on ImageSet2Text’s value for visually-impaired individuals, incorporating community insights early in the design process (costanza2020). Details in App. F.
@@ 6 Conclusion
In this paper, we have proposed ImageSet2Text, a novel method to generate natural language descriptions of image sets. We have shown its competitive performance in: (1) a large-scale group captioning experiment with two newly proposed benchmarks (GroupConceptualCaptions and GroupWikiArt); (2) set difference captioning to assess description completeness; and (3) a user study. Through an ablation study, a scalability analysis, and failure case examination, we have illustrated how ImageSet2Text successfully integrates symbolic and data-centric approaches. Overall, ImageSet2Text effectively describes large image collections, proving valuable for diverse applications.
@@ Acknowledgements
We are grateful to Fundación ONCE for their willingness to collaborate on our research. PR and NO have been partially supported by a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Resolución de la Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación). PR was also supported by a grant by Fundación Banc Sabadell. NG is partly supported by JSPS KAKENHI No. JP22K12091. KS is part of the ELLIS Unit Linz, the LIT AI Lab and the Institute for Machine Learning at Johannes Kepler University Linz, which are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG. We thank Federico Brunero, Julien Colin, Erik Derner, Aditya Gulati, Benedikt Höltgen, Fabian Paischer and Korbinian Pöppel for helpful discussions.
Appendix A ImageSet2Text: Implementation Details
@@
To effectively prompt GPT-4o-mini, we design system prompts that outline the model’s role and guide its reasoning process. Additionally, we utilize structured outputs to ensure consistency in responses. For example, we reference the “next question” prompt used in ImageSet2Text, provided in Fig. 7, to explain the rationale behind our prompt engineering. This is a key component in our design, as it provides a structured method for generating follow-up questions to refine image set descriptions.
The system prompt defines GPT-4o-mini’s role as an expert assisting in generating textual descriptions for a large image set. Specifically, the model is given:
@@
KEY_POINT: The subject at the current iteration, i.e., the main element within the description that needs additional details.
ATTRIBUTE: The predicate at the current iteration, i.e., the specific property of the KEY_POINT to investigate further.
@@
This structured approach overall improves the quality of GPT-4o-mini. The other prompts used for the LLM in ImageSet2Text follow the same criteria. As a further example, in Figure 8 we also provide the graph-to-text rendering prompt, used at the final stage of the pipeline of ImageSet2Text.
1NEXT_QUESTION = """You are an expert in {expertise} assisting a client in enriching the textual description of a large IMAGE_SET in their possession.
2
3Using JSON, you will be provided with a question expressed as a connection within a knowledge graph built for IMAGE_SET (QUESTION_BRANCH), a KEY_POINT that lacks detailed explanation in DESCRIPTION_CURRENT, an ATTRIBUTE which is the specific property of the KEY_POINT to investigate further, and a list of textual questions which have already been asked about that KEY_POINT (LOG).
4
5An example of a client query might be the following:
6
7{jsonScheme_input}
8
9Based on this input, translate the QUESTION_BRANCH into two textual concise questions (QUESTION_EXPERT and QUESTION_VQA) to investigate the ATTRIBUTE further.
10
11Provide the answer in JSON format. For example, the output from an expert in couples would be:
12
13{jsonScheme_output}
14
15In the output, it is expected that:
16- The field QUESTION_EXPERT is a string containing an expert-level query that specifically addresses the selected ATTRIBUTE, intended to enrich the understanding of the IMAGE_SET.
17- The field QUESTION_VQA is a string containing the translation of QUESTION_EXPERT into an image-focused question that meets the following criteria:
18 1. QUESTION_VQA must be direct, simple, clear, and must refer to visible aspects within the image.
19 2. QUESTION_VQA should be structured to ask about details that can be observed in one randomly selected image from the IMAGE_SET, referring to "the image" directly, like "What is in the image?".
20 3. Avoid questions that require abstract reasoning, cultural knowledge, or expertise beyond what can be visually interpreted.
21 4. Avoid yes/no questions and focus instead on generating descriptive responses.
22 5. Ensure that QUESTION_VQA aligns with the capabilities of a VQA system, such as object recognition, spatial relationships, counting, and scene understanding.
23 6. QUESTION_VQA must be designed to elicit new insights without overlapping with previous responses.
@@
1def next_question(expertise):
2 expertise = expertise_to_string(expertise)
3
4 jsonScheme_input = """{
5 "QUESTION_BRANCH": "image.wedding.couple.body language?",
6 "KEY_POINT": "couple",
7 "ATTRIBUTE": "body language: communication via the movements or attitudes of the body",
8 "log": [
9 {
10 "ATTRIBUTE": "gender composition: the properties that distinguish organisms on the basis of their reproductive roles",
11 "QUESTION_EXPERT": "What is the gender composition of the couple portrayed in the image?",
12 "QUESTION_VQA": "How many men and women are visible in the image?"
13 },
14 {
15 "ATTRIBUTE": "attire: clothing of a distinctive style or for a particular occasion",
16 "QUESTION_EXPERT": "What is the attire of the couple portrayed in the image?",
17 "QUESTION_VQA": "What type of clothing are the individuals in the image wearing?"
18 }
19 ]
20 }"""
21
22 jsonScheme_output = """{
23 "QUESTION_EXPERT": "How does the couple’s body language appear in the wedding image?",
24 "QUESTION_VQA": "What are the body positions or movements of the spouses in the image?"
25 }"""
26
27 return NEXT_QUESTION.format(
28 expertise=expertise,
29 jsonScheme_input=jsonScheme_input,
30 jsonScheme_output=jsonScheme_output
31 )’
@@
1WRITE_DESCRIPTION = """You are assisting a client in writing the textual description of a large IMAGE_SET in their possession.
2
3The client has collected pieces of verified information about the IMAGE_SET in form of a knowledge graph, and has generated a network text representation out of it.
4
5Using JSON, you will be provided with the knowledge graph text representation (GRAPH).
6
7An example of a client query might be the following:
8
9{jsonScheme_input}
10
11Given the GRAPH, write a textual description (DESCRIPTION) of the entire IMAGE_SET. Provide the answer in JSON format. For example:
12
13{jsonScheme_output}
14
15GUIDELINES
16- As you write the DESCRIPTION, make sure to retain all existing details from the GRAPH. Do not miss any piece of information.
17- Stick strictly to the details provided in the GRAPH without adding, assuming, or inventing any information.
18- Note that all information found in the GRAPH applies to every image within the IMAGE_SET.
19"""
(a) “Graph to Text” Prompt
1def write_description():
2 jsonScheme_input = " {" + """
3 "GRAPH": "-- (’entity’, ’image’)
4 |-> (’image’, ’content’, ’mountains’)
5 | |-> (’mountains’, ’portrayed_during’, ’golden hour’)
6 |-> (’image’, ’type’, ’photographs’)"
7 """ + "}"
8
9 jsonScheme_output = " {" + """
10 "DESCRIPTION": "A collection of photographs capturing mountains, portrayed during the golden hour."
11 """ + "}"
12
13 return WRITE_DESCRIPTION.format(jsonScheme_input=jsonScheme_input, jsonScheme_output=jsonScheme_output)
(b) Function for “Next Question”
Figure 8: “Graph 2 Text” Prompt
Appendix B Accuracy Evaluation
In this section, we provide additional information regarding the accuracy evaluation. Specifically, we provide details about the dataset construction process, how to generate a caption with ImageSet2Text, details on how the baselines have been set up and the full results of the experiments.
B.1 Dataset Composition and Construction